• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Subedi

    Log entry time 11:40:01 on January21,2009

    Entry number 257119

    This entry is a followup to: 257102

    keyword=BPM lock for Compton

    During the beam-study yesterday, the lock for IPM1P02B was tunrned off and it remained off (had to be turned ON) until today (Jan 21) 08:20. This was the reason for not seeing the Compton singal and background separated. As soon as the BPM lock was turned on, the Compton signal and background appeared separated. Until now, the Compton signal is very good.